# Constraint Learning for Parametric Point Cloud

### v3

@@ C.2 Dataset Statics
Shapes instantiated from the parametric templates designed by our team names begin with "Num".
@@
Figure 13: Quantity distribution of Parametric 20000 dataset.
Figure 14: Shapes proportion of Parametric 20000 dataset.
Figure 15: Example shapes for each category in Parametric 20000 dataset.
D Constraint Learning Method Details
D.1 Defects of General Constraint Representation
@@ 13                   Return $v$;
D.3 CST-BRep Module
This section explains why it is necessary to compute valid edges in the CST-BRep module.
The CST-BRep module is used to compute constraints from B-Rep data, and the computation process consists of the following steps:
@@ D.3 CST-BRep Module
D.4 CST-PCD Module
This section provides a detailed explanation of the valid points in the computational steps that inspire the design of the CST-PCD Module.
The CST-PCD Module is used to predict constraints from point clouds. The design of the CST-PCD Module is inspired by the following calculation process:
@@ D.5 SurfaceKNN
E Classification Segmentation Experiments
E.1 Process to Visualize Main Axis Direction
E.2 Constraint Visualization of CST-BRep
More constraints extracted by the CST-BRep module from the B-Rep are visualized in Fig. 22.
Figure 22: More constraint extraction results from the B-Rep data. ADJ: Adjacency, PT: Primitive Type, MAD: Main Axis Direction.
E.3 Segmentation
The full segmentation results of investigated methods on 360 gallery are shown in Tab. 6.
Table 6: Part Segmentation Results on 360 Gallery. OA: overall accuracy %. mIoU: instance mIoU %. mIoU*: class mIoU %. Segmentation categories from left to right are: Extrude Side, Extrude End, Cut Side, Cut End, Fillet, Chamfer, Revolve Side, Revolve End. PN: PointNet [33]. PN2: PointNet++ [34]. DGN: DGCNN [42]. 3DN: 3DGCN [26].
Method PN PN2 DGN 3DN Ours
OA 42.13 58.18 64.02 60.58 70.76
mIoU 18.89 26.98 35.54 33.03 46.65
mIoU* 7.04 11.27 14.36 12.23 20.51
[Uncaptioned image] 19.83 49.80 54.48 50.67 61.29
[Uncaptioned image] 36.47 40.36 50.82 47.20 60.65
[Uncaptioned image] 0 0 0 0 0
[Uncaptioned image] 0 0 0 0 0
[Uncaptioned image] 0 0 0 0 0
[Uncaptioned image] 0 0 0 0 0
[Uncaptioned image] 0 0 9.60 0 0
[Uncaptioned image] 0 0 0 0 0

### v5

@@ H.3 Dataset Statics
Shapes instantiated from the parametric templates designed by our team names begin with ”Num”.
@@
Figure 14: Quantity distribution of Param20K dataset.
Figure 15: Shapes proportion of Param20K dataset.
Figure 16: Example shapes for each category in Param20K dataset.
I Constraint Acquisition and Classification Experiments
I.1 Process to Visualize Main Axis Direction
I.2 Constraint Visualization of CstBRep
More constraints extracted by the CstBRep module from the B-Rep are visualized in Fig. 17.
Figure 17: More constraint extraction results from the B-Rep data. ADJ: Adjacency, PMT: Primitive Type, MAD: Main Axis Direction.
I.3 Constraint acquisition on MCB
We evaluated the pre-trained CstPnt module and ParSeNet [41], HPNet [58] on the CAD shape dataset MCB [18] to predict constraints. For the MCB consists of mesh files, it was not possible to extract Ground Truth constraints for comparison, therefore no ground truth presented in Fig. 18, and no numerical results are available. Based on the results visualized in Fig. 18, CstPnt demonstrated a high level of constraint prediction accuracy even on an unseen dataset.
Figure 18: Constraint prediction on MCB.
I.4 Classification on MCB
We conducted evaluations on the CAD shape dataset MCB [18], with results presented in Tab. 5, from which our CstNet achieved the highest scores across all metrics, conclusions similar with experiments on Param20K dataset.
Table 5: Classification results on MCB. Acc: accuracy over instance %, Acc*: accuracy over class %, F1: F1-score, mAP: mean average precision %.
Method Acc Acc* F1 mAP
PointCNN [27] 93.89 81.85 83.86 90.13
PointNet [36] 86.78 67.70 86.55 74.08
PointNet++ [37] 87.45 73.68 88.32 91.33
SpiderCNN [57] 93.59 79.70 81.30 86.64
PointConv [51] 93.25 80.24 71.31 82.19
DGCNN [46] 92.54 74.47 76.12 74.27
3DGCN [29] 93.71 78.71 84.59 84.35
Ours 96.87 89.21 89.85 93.17
J Constraint Learning Method Details
J.1 Defects of General Constraint Representation
@@ 13                   Return $v$;
J.3 CstBRep Module
This section explains why it is necessary to compute valid edges in the CstBRep module.
The CstBRep module is used to compute constraints from B-Rep data, and the computation process consists of the following steps:
@@
J.4 CstPnt Module
This section provides a detailed explanation of the valid points in the computational steps that inspire the design of the CstPnt Module.
The CstPnt Module is used to predict constraints from point clouds. The design of the CstPnt Module is inspired by the following calculation process: